# Structural Alignment Improves Graph Test-Time Adaptation

### v1

@@ B.2 Results of Ablation Study
TENT+TSA Base 27.30$\pm$1.61 32.84$\pm$0.78 40.82$\pm$0.99 32.53$\pm$0.91 28.62$\pm$0.63 27.89$\pm$0.98 16.22$\pm$1.17 24.87$\pm$0.56 22.05$\pm$0.84 14.20$\pm$0.71
w/o Nbr. Align 27.18$\pm$1.70 32.75$\pm$0.69 40.82$\pm$0.99 32.26$\pm$0.95 28.41$\pm$0.59 27.53$\pm$0.92 16.22$\pm$1.17 24.87$\pm$0.56 22.05$\pm$0.84 13.95$\pm$0.76
w/o SNR 27.00$\pm$1.21 32.76$\pm$0.75 40.80$\pm$0.91 32.52$\pm$0.90 28.57$\pm$0.59 27.70$\pm$0.88 15.84$\pm$0.85 24.62$\pm$0.48 21.37$\pm$0.73 14.11$\pm$0.60
LAME+TSA Base 37.95$\pm$0.97 36.29$\pm$1.65 46.86$\pm$1.13 32.86$\pm$2.23 27.22$\pm$1.48 44.83$\pm$0.88 28.51$\pm$0.44 39.80$\pm$0.99 24.54$\pm$0.87 22.39$\pm$0.30
w/o Nbr. Align 37.90$\pm$0.97 36.08$\pm$1.73 46.75$\pm$0.88 32.70$\pm$2.29 26.49$\pm$1.76 44.83$\pm$0.88 28.41$\pm$0.36 39.72$\pm$0.99 24.24$\pm$0.75 22.23$\pm$0.32
w/o SNR 35.35$\pm$0.84 33.45$\pm$1.60 44.51$\pm$1.22 29.89$\pm$1.94 24.41$\pm$1.43 40.17$\pm$1.11 22.44$\pm$0.99 32.57$\pm$1.46 17.33$\pm$0.47 17.07$\pm$0.61
T3A+TSA Base 41.65$\pm$0.99 47.01$\pm$2.08 51.65$\pm$0.90 46.61$\pm$0.88 43.45$\pm$0.81 48.09$\pm$0.60 39.18$\pm$1.87 46.50$\pm$0.25 43.70$\pm$1.38 30.89$\pm$2.13
w/o Nbr. Align 41.64$\pm$1.10 46.37$\pm$2.11 50.32$\pm$1.11 46.44$\pm$0.80 42.09$\pm$1.26 46.47$\pm$1.27 39.12$\pm$1.85 46.23$\pm$0.61 43.60$\pm$1.21 30.92$\pm$2.03
w/o SNR 41.59$\pm$0.94 45.54$\pm$1.81 51.55$\pm$0.95 46.51$\pm$0.94 41.35$\pm$1.09 48.11$\pm$0.62 38.71$\pm$1.00 46.29$\pm$0.26 43.12$\pm$0.76 29.90$\pm$1.35
B.3 Visualization of Distribution Shift
(a) Source (b) Feature Shift (c) Neighborhood Shift (d) Label Shift (e) SNR
Figure 4: The t-SNE visualization of a one-layer GSN under different distribution shifts: (top) the mean-pooled neighborhood hidden representations and (bottom) the node representations before passing through the classifier, positioned along its decision boundary. (a) indicates the source domain where the model is pretrained. (b), (c), and (d) stand for feature shift, neighborhood shift, and label shift. (e) represent the impact of SNR induced by degree shift. The color of the nodes represents the ground-truth labels. The legends in the bottom plots show the accuracy and recall scores for each class. The red contours in the bottom (d) highlight the decision boundary of the minority class.
Appendix C Dataset Details
C.1 Dataset Statistics
@@
Here we summarize the statistics of the four real-world graph datasets. The edges are counted twice in the edge list as they are undirected graphs.
@@
gg-10 qq-10 gg-30 qq-30 gg-50 gg-140
Nodes $18611$ $17242$ $41390$ $38929$ $60054$ $154750$
Edges $53725$ $42769$ $173392$ $150026$ $341930$ $2081229$
Features $28$ $28$ $28$ $28$ $28$ $28$
Classes $2$ $2$ $2$ $2$ $2$ $2$
C.2 Dataset Shift Statistics
Below we provide two metrics indicating the neighborhood and label shift in real-world dataset. We measure the neighborhood shift by computing the weighted average of the total variation distance between of the neighborhood distribution:
[…]
@@ C.2 Dataset Shift Statistics
[…]
@@
$US\rightarrow CN$ $US\rightarrow DE$ $US\rightarrow JP$ $US\rightarrow RU$ $US\rightarrow FR$ $CN\rightarrow US$ $CN\rightarrow DE$ $CN\rightarrow JP$ $CN\rightarrow RU$ $CN\rightarrow FR$ Nbr. Shift $0.2137$ $0.1902$ $0.1476$ $0.2855$ $0.2238$ $0.1946$ $0.2053$ $0.1361$ $0.2025$ $0.2039$ Label Shift $0.2734$ $0.1498$ $0.1699$ $0.3856$ $0.1706$ $0.2734$ $0.2691$ $0.1522$ $0.2453$ $0.2256$
@@
Pileup Conditions Physical Processes Domains $\text{PU}10\rightarrow 30$ $\text{PU}30\rightarrow 10$ PU1050105010$\rightarrow$ 5010 → 50 PU5010501050$\rightarrow$ 1050 → 10 PU301403014030$\rightarrow$ 14030 → 140 PU1403014030140$\rightarrow$ 30140 → 30 $gg\rightarrow qq$ $qq\rightarrow gg$ Nbr. Shift $0.1493$ $0.1709$ $0.2039$ $0.2595$ $0.1293$ $0.1635$ $0.0214$ $0.0205$ Label Shift $0.2425$ $0.2425$ $0.2982$ $0.2982$ $0.1377$ $0.1377$ $0.0341$ $0.0348$
@@
1950-2007 1950-2009 1950-2011 DBLP/ACM Domains $2014-2016$ $2016-2018$ $2014-2016$ $2016-2018$ $2014-2016$ $2016-2018$ $D\rightarrow A$ $A\rightarrow D$ Nbr. Shift $0.3005$ $0.3450$ $0.2583$ $0.3120$ $0.1833$ $0.2567$ $0.1573$ $0.2227$ Label Shift $0.2938$ $0.4396$ $0.2990$ $0.4552$ $0.2853$ $0.4438$ $0.3434$ $0.3435$
Appendix D Detailed Experimental Setup
D.1 Datasets Setup
For Arxiv, MAG, and DBLP/ACM we follow the setting of (Liu et al., 2024). Hence we discuss the setting of synthetic dataset and pileup in the following:

### v2

@@ D.4 Hyperparameter sensitivity
ERM 50.60$\pm$3.52 68.09$\pm$0.74 32.60$\pm$5.55 64.59$\pm$1.16 0.68$\pm$0.40 24.95$\pm$5.32 71.00$\pm$0.61 69.56$\pm$1.06
GTrans 51.08$\pm$3.06 66.26$\pm$1.77 33.94$\pm$5.39 62.53$\pm$1.11 0.75$\pm$0.52 25.46$\pm$5.19 70.11$\pm$1.13 68.40$\pm$1.93
HomoTTT 52.05$\pm$3.09 67.46$\pm$0.94 35.84$\pm$4.98 63.59$\pm$1.11 0.75$\pm$0.43 23.03$\pm$5.46 71.06$\pm$0.50 69.72$\pm$1.10
SOGA 54.18$\pm$3.08 66.13$\pm$0.70 42.30$\pm$4.74 59.97$\pm$1.70 1.16$\pm$0.84 18.68$\pm$5.49 69.72$\pm$1.27 68.39$\pm$1.75
ActMAD 58.80$\pm$1.58 62.22$\pm$0.95 49.22$\pm$2.21 56.04$\pm$2.24 2.16$\pm$0.75 17.81$\pm$6.45 70.40$\pm$0.94 69.95$\pm$1.02
TENT 36.64$\pm$0.15 2.04$\pm$1.05 27.01$\pm$0.06 0.62$\pm$0.81 12.74$\pm$0.50 3.77$\pm$3.30 60.19$\pm$3.55 71.19$\pm$0.58
LAME 46.33$\pm$4.67 67.61$\pm$0.96 22.58$\pm$9.42 64.58$\pm$1.00 0.00$\pm$0.00 20.94$\pm$7.27 70.72$\pm$0.74 69.67$\pm$0.88
T3A 55.66$\pm$2.89 71.36$\pm$0.79 47.89$\pm$4.19 71.01$\pm$0.38 19.69$\pm$3.72 54.38$\pm$2.56 70.34$\pm$1.33 68.63$\pm$1.51
Matcha-TENT 36.58$\pm$0.17 0.00$\pm$0.00 27.01$\pm$0.06 0.00$\pm$0.00 9.17$\pm$0.94 0.00$\pm$0.00 0.72$\pm$0.76 20.30$\pm$4.03
Matcha-LAME 7.45$\pm$3.55 0.00$\pm$0.00 8.28$\pm$4.16 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 0.52$\pm$1.05 6.56$\pm$4.08
Matcha-T3A 15.48$\pm$2.91 0.00$\pm$0.00 14.15$\pm$1.86 0.00$\pm$0.00 0.00$\pm$0.00 9.60$\pm$19.19 17.20$\pm$14.72 22.77$\pm$5.23
TSA-TENT 39.90$\pm$1.18 5.07$\pm$10.13 27.78$\pm$0.48 0.59$\pm$0.75 20.31$\pm$0.65 3.65$\pm$3.22 67.97$\pm$1.60 71.60$\pm$0.18
TSA-LAME 60.33$\pm$0.80 71.54$\pm$0.33 52.23$\pm$0.61 70.62$\pm$0.35 38.32$\pm$0.27 20.49$\pm$7.76 72.28$\pm$0.47 71.23$\pm$0.27
TSA-T3A 59.40$\pm$1.70 72.94$\pm$0.18 53.13$\pm$0.34 72.63$\pm$0.23 36.39$\pm$1.48 62.31$\pm$0.28 71.30$\pm$1.23 69.70$\pm$1.04
Table 9: GPRGNN results on Arxiv and DBLP/ACM (accuracy). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.
1950-2007 1950-2009 1950-2011 DBLP & ACM
[…]
ERM 45.62$\pm$2.30 40.35$\pm$4.46 49.21$\pm$1.36 44.78$\pm$2.70 56.49$\pm$0.97 54.14$\pm$1.31 67.34$\pm$1.21 56.47$\pm$2.84
GTrans 45.25$\pm$2.25 40.51$\pm$4.31 48.56$\pm$1.23 44.57$\pm$2.55 55.81$\pm$0.99 53.59$\pm$1.34 59.61$\pm$0.98 59.27$\pm$1.79
HomoTTT 45.49$\pm$2.64 40.03$\pm$4.80 48.10$\pm$1.17 42.33$\pm$2.34 56.27$\pm$1.03 53.25$\pm$1.24 67.11$\pm$1.03 57.18$\pm$2.54
SOGA 48.40$\pm$2.29 46.01$\pm$4.00 52.59$\pm$1.38 50.64$\pm$1.63 58.61$\pm$1.15 57.68$\pm$0.88 67.00$\pm$1.03 55.42$\pm$3.03
ActMAD 45.47$\pm$2.28 40.12$\pm$4.42 49.42$\pm$1.31 44.96$\pm$2.59 56.63$\pm$1.00 54.25$\pm$1.25 67.07$\pm$1.11 55.83$\pm$2.76
TENT 44.73$\pm$2.63 40.58$\pm$3.44 48.47$\pm$0.94 42.37$\pm$1.52 57.15$\pm$0.98 52.45$\pm$1.75 61.56$\pm$0.77 56.84$\pm$1.04
LAME 45.75$\pm$3.96 41.35$\pm$7.20 49.52$\pm$2.24 50.18$\pm$3.36 55.65$\pm$1.09 54.93$\pm$1.72 66.92$\pm$0.99 56.58$\pm$2.83
T3A 32.73$\pm$1.63 31.97$\pm$1.75 36.99$\pm$2.01 34.10$\pm$1.32 40.53$\pm$2.88 39.04$\pm$3.86 67.21$\pm$1.31 65.32$\pm$1.69
Matcha-TENT 43.89$\pm$2.77 39.30$\pm$3.63 48.01$\pm$0.90 41.91$\pm$1.66 56.28$\pm$1.30 51.27$\pm$1.64 64.43$\pm$0.90 78.61$\pm$1.38
Matcha-LAME 32.93$\pm$3.30 24.13$\pm$5.06 37.41$\pm$2.06 24.16$\pm$2.92 48.29$\pm$3.27 44.74$\pm$4.20 67.83$\pm$1.71 70.39$\pm$2.15
Matcha-T3A 24.81$\pm$2.32 19.06$\pm$3.09 25.85$\pm$3.54 19.60$\pm$4.56 26.90$\pm$1.55 18.74$\pm$4.37 67.59$\pm$1.87 72.91$\pm$2.69
TSA-TENT 45.17$\pm$2.84 41.07$\pm$3.67 49.13$\pm$0.84 43.19$\pm$1.24 57.66$\pm$1.00 53.00$\pm$1.65 62.77$\pm$0.84 61.71$\pm$0.99
TSA-LAME 49.75$\pm$4.12 48.67$\pm$7.74 53.82$\pm$0.53 55.18$\pm$1.21 57.67$\pm$0.98 57.14$\pm$1.83 67.71$\pm$1.10 65.20$\pm$1.70
TSA-T3A 39.11$\pm$1.73 39.40$\pm$3.69 41.15$\pm$6.33 44.65$\pm$2.31 44.14$\pm$2.63 40.82$\pm$5.28 68.02$\pm$1.40 74.64$\pm$1.38
Table 10: GCN results on MAG (accuracy). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.
[…]
ERM 27.76$\pm$1.49 30.70$\pm$0.96 40.12$\pm$0.73 27.70$\pm$0.75 23.03$\pm$0.80 36.46$\pm$2.00 21.03$\pm$1.56 28.95$\pm$2.43 20.53$\pm$2.07 16.68$\pm$1.54
GTrans 27.44$\pm$1.48 30.59$\pm$0.99 39.65$\pm$0.97 27.79$\pm$0.72 22.95$\pm$0.88 35.88$\pm$2.09 20.80$\pm$1.59 28.13$\pm$2.49 20.53$\pm$2.06 16.52$\pm$1.51
HomoTTT 23.64$\pm$2.09 28.49$\pm$1.29 35.34$\pm$1.87 26.21$\pm$1.62 23.78$\pm$0.56 29.27$\pm$2.91 18.60$\pm$1.59 23.10$\pm$2.01 21.71$\pm$1.70 16.03$\pm$1.33
SOGA 28.36$\pm$2.23 35.61$\pm$1.50 45.29$\pm$2.08 36.79$\pm$1.18 29.30$\pm$0.90 44.09$\pm$1.92 26.78$\pm$1.26 37.93$\pm$2.05 33.24$\pm$1.49 24.23$\pm$1.01
ActMAD 28.11$\pm$1.34 31.25$\pm$0.90 40.71$\pm$0.65 28.61$\pm$0.63 23.82$\pm$0.74 36.07$\pm$2.01 20.73$\pm$1.50 28.76$\pm$2.33 20.92$\pm$2.04 16.65$\pm$1.46
TENT 26.66$\pm$0.57 33.38$\pm$0.32 41.84$\pm$0.59 33.01$\pm$0.54 29.17$\pm$0.20 28.82$\pm$0.85 16.35$\pm$0.76 25.20$\pm$0.85 23.45$\pm$0.87 14.96$\pm$0.56
LAME 28.20$\pm$3.24 30.90$\pm$1.77 42.25$\pm$0.97 27.20$\pm$2.17 20.83$\pm$1.08 35.89$\pm$2.53 19.31$\pm$2.11 30.01$\pm$4.10 17.46$\pm$2.39 14.24$\pm$2.02
T3A 41.53$\pm$0.93 45.41$\pm$1.76 51.11$\pm$1.18 48.02$\pm$1.99 40.09$\pm$1.53 43.72$\pm$2.31 39.68$\pm$2.25 47.31$\pm$0.84 45.93$\pm$1.14 30.95$\pm$1.20
TSA-TENT 27.46$\pm$0.53 33.89$\pm$0.31 42.34$\pm$0.60 33.72$\pm$0.53 29.59$\pm$0.21 29.33$\pm$0.86 16.82$\pm$0.75 25.90$\pm$0.89 24.03$\pm$0.81 15.32$\pm$0.56
TSA-LAME 28.64$\pm$3.24 31.01$\pm$1.91 42.35$\pm$0.86 27.22$\pm$1.87 22.27$\pm$0.52 39.64$\pm$2.08 22.96$\pm$1.62 33.94$\pm$3.14 20.08$\pm$1.93 17.36$\pm$1.91
TSA-T3A 41.61$\pm$0.95 46.13$\pm$1.83 51.81$\pm$1.22 48.87$\pm$1.43 42.98$\pm$1.90 45.48$\pm$2.69 40.11$\pm$2.74 47.45$\pm$0.42 45.95$\pm$1.17 32.03$\pm$1.30
Table 11: GCN results on Pileup (f1-scores). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.
[…]
ERM 8.30$\pm$1.79 29.20$\pm$4.57 2.27$\pm$0.70 26.61$\pm$6.79 0.33$\pm$0.13 16.08$\pm$10.46 23.71$\pm$2.49 26.73$\pm$4.79
GTrans 11.97$\pm$2.67 28.88$\pm$4.71 3.90$\pm$1.26 27.20$\pm$6.88 0.91$\pm$0.29 16.98$\pm$11.06 24.62$\pm$2.38 28.88$\pm$5.14
HomoTTT 8.94$\pm$2.05 31.21$\pm$4.79 2.54$\pm$0.81 27.59$\pm$6.18 0.55$\pm$0.18 16.77$\pm$11.23 24.84$\pm$3.27 27.48$\pm$4.95
SOGA 5.81$\pm$2.20 35.21$\pm$4.00 1.00$\pm$0.44 27.19$\pm$6.76 0.54$\pm$0.20 19.17$\pm$11.32 25.38$\pm$1.66 28.06$\pm$5.06
ActMAD 7.20$\pm$2.22 30.45$\pm$4.13 1.15$\pm$0.41 24.35$\pm$6.69 0.31$\pm$0.14 16.67$\pm$9.32 23.77$\pm$1.78 27.36$\pm$4.75
TENT 2.45$\pm$0.93 35.99$\pm$3.12 1.63$\pm$0.75 26.20$\pm$7.51 0.09$\pm$0.05 13.84$\pm$8.89 19.96$\pm$4.06 29.58$\pm$4.46
LAME 7.48$\pm$2.17 30.91$\pm$5.58 1.46$\pm$0.75 24.97$\pm$9.02 0.06$\pm$0.06 15.30$\pm$10.74 21.50$\pm$5.24 25.72$\pm$6.51
T3A 41.55$\pm$13.17 62.80$\pm$2.48 29.92$\pm$7.29 56.32$\pm$14.68 4.45$\pm$2.48 33.26$\pm$4.35 60.15$\pm$10.72 62.77$\pm$8.69
TSA-TENT 2.96$\pm$1.12 38.45$\pm$3.83 2.17$\pm$0.82 26.85$\pm$7.89 0.24$\pm$0.16 15.20$\pm$10.17 24.22$\pm$3.88 33.47$\pm$4.43
TSA-LAME 12.71$\pm$2.29 32.37$\pm$8.02 4.14$\pm$1.48 27.14$\pm$10.60 0.38$\pm$0.19 16.53$\pm$12.07 29.29$\pm$2.32 27.72$\pm$4.22
TSA-T3A 43.86$\pm$13.14 64.00$\pm$1.56 33.31$\pm$8.50 60.22$\pm$8.38 7.12$\pm$4.19 35.72$\pm$4.56 62.03$\pm$8.01 64.02$\pm$6.87
Table 12: GCN results on Arxiv and DBLP/ACM (accuracy). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.
1950-2007 1950-2009 1950-2011 DBLP & ACM
[…]
ERM 49.25$\pm$1.58 50.64$\pm$3.11 52.61$\pm$2.67 51.92$\pm$4.12 58.83$\pm$0.83 57.73$\pm$1.27 62.56$\pm$3.32 47.59$\pm$1.92
GTrans 49.10$\pm$1.68 50.57$\pm$3.10 52.50$\pm$2.71 51.87$\pm$4.26 58.72$\pm$0.89 57.75$\pm$1.31 62.03$\pm$2.31 51.65$\pm$3.09
HomoTTT 46.22$\pm$1.54 46.11$\pm$2.79 48.96$\pm$3.13 46.83$\pm$5.00 55.10$\pm$1.63 53.18$\pm$2.52 63.21$\pm$3.31 48.45$\pm$1.68
SOGA 49.82$\pm$2.38 52.36$\pm$2.54 52.92$\pm$1.08 54.45$\pm$1.15 57.93$\pm$1.71 58.66$\pm$2.02 64.22$\pm$2.40 53.29$\pm$2.08
ActMAD 49.24$\pm$1.57 50.68$\pm$2.94 52.71$\pm$2.64 52.10$\pm$3.98 58.83$\pm$0.86 57.85$\pm$1.25 62.68$\pm$3.36 47.58$\pm$1.89
TENT 48.59$\pm$1.41 49.90$\pm$1.15 53.39$\pm$1.82 53.45$\pm$2.11 58.72$\pm$0.91 58.23$\pm$0.95 60.80$\pm$2.51 48.11$\pm$1.43
LAME 49.77$\pm$2.50 50.97$\pm$4.60 53.35$\pm$2.72 52.70$\pm$4.25 58.85$\pm$0.57 57.05$\pm$1.57 62.51$\pm$3.33 47.64$\pm$1.70
T3A 43.37$\pm$1.38 41.32$\pm$3.13 49.95$\pm$3.76 47.11$\pm$6.73 55.56$\pm$1.16 53.57$\pm$1.94 62.60$\pm$3.65 49.65$\pm$1.83
TSA-TENT 49.63$\pm$1.26 50.78$\pm$0.74 53.78$\pm$1.57 53.67$\pm$1.85 58.90$\pm$0.94 58.27$\pm$0.95 61.31$\pm$2.39 53.14$\pm$2.01
TSA-LAME 50.95$\pm$2.05 53.04$\pm$2.86 54.49$\pm$1.81 54.13$\pm$3.35 59.26$\pm$0.70 57.55$\pm$1.30 63.92$\pm$2.76 55.19$\pm$4.50
TSA-T3A 45.37$\pm$0.76 45.46$\pm$1.92 51.57$\pm$2.73 49.70$\pm$5.44 56.14$\pm$1.02 54.84$\pm$1.17 63.31$\pm$3.50 54.53$\pm$3.07
D.5 Results of More GNN Architectures
TSA is Model-Agonistic.  TSA employs neighborhood alignment and SNR adjustment on edge weights and do not need any assumption on the GNN architecture such as hop-aggregation parameters. We demonstrate that  TSA is applicable to common GNN architectures such as GraphSAGE, GPRGNN, and GCN. While different backbones yield varying ERM results, we observe that  TSA consistently improves model performance.
Comparison with Baselines. TSA exhibits strong outcomes on real-world datasets and with both GPRGNN and GCN backbones. In all cases, the best results are obtained by GTTA methods (underlined across Table712), highlighting the importance of addressing structure shift in graph data. For GPRGNN,  TSA consistently outperforms GTrans and HomoTTT and surpasses SOGA by an average of 10.5% and Matcha by an average of 19.5%. For GCN,  TSA similarly outperforms GTrans and HomoTTT, and exceeds SOGA by an average of 13.1%. Furthermore, TSA shows prevailing gains over the corresponding non-graph TTA baselines (bold across Table712). While Matcha sometimes demonstrates strong performance, we frequently observe unstable results that may even fall below ERM. This drop is attributed to the reliance of their proposed PIC loss on minimizing intra-variance, which depends on the quality of node hidden representations in the target domain. Under label shift, node representations can easily degrade, leading to instability in Matcha’s performance.
Appendix E Dataset Details
E.1 Dataset Statistics
@@
Here we summarize the statistics of the four real-world graph datasets in Table LABEL:table:arxiv_datastats, LABEL:table:mag_datastats, and LABEL:table:pileup_datastats. The edges are counted twice in the edge list as they are undirected graphs.
@@
gg-10 qq-10 gg-30 gg-50 gg-140
Nodes $18611$ $17242$ $41390$ $60054$ $154750$
Edges $53725$ $42769$ $173392$ $341930$ $2081229$
Features $28$ $28$ $28$ $28$ $28$
Classes $2$ $2$ $2$ $2$ $2$
E.2 Dataset Shift Statistics
Below we provide two metrics indicating the neighborhood and label shift in real-world dataset. We measure the CSS by computing the weighted average of the total variation distance between of the neighborhood distribution:
$\text{CSS}\coloneqq\frac{1}{2}\sum_{j\in{\mathcal{Y}}}\sum_{k\in{\mathcal{Y}}}\mathbb{P}^{\mathcal{T}}(Y_{u}=j)\lvert\mathbb{P}^{\mathcal{S}}(Y_{v}=k|Y_{u}=j,v\in\mathcal{N}_{u})-\mathbb{P}^{\mathcal{T}}(Y_{v}=k|Y_{u}=j,v\in\mathcal{N}_{u})\rvert$
@@ E.2 Dataset Shift Statistics
Label $\text{Label Shift}\coloneqq\frac{1}{2}\sum_{j\in{\mathcal{Y}}}\lvert\mathbb{P}^{\mathcal{S}}(Y_{u}=j)-\mathbb{P}^{\mathcal{T}}(Y_{u}=j)\rvert$
@@
$US\rightarrow CN$ $US\rightarrow DE$ $US\rightarrow JP$ $US\rightarrow RU$ $US\rightarrow FR$ $CN\rightarrow US$ $CN\rightarrow DE$ $CN\rightarrow JP$ $CN\rightarrow RU$ $CN\rightarrow FR$ CSS $0.2137$ $0.1902$ $0.1476$ $0.2855$ $0.2238$ $0.1946$ $0.2053$ $0.1361$ $0.2025$ $0.2039$ Label Shift $0.2734$ $0.1498$ $0.1699$ $0.3856$ $0.1706$ $0.2734$ $0.2691$ $0.1522$ $0.2453$ $0.2256$
@@
Pileup Conditions Physical Processes Domains $\text{PU}10\rightarrow 30$ $\text{PU}30\rightarrow 10$ PU1050105010$\rightarrow$ 5010 → 50 PU5010501050$\rightarrow$ 1050 → 10 PU301403014030$\rightarrow$ 14030 → 140 PU1403014030140$\rightarrow$ 30140 → 30 $gg\rightarrow qq$ $qq\rightarrow gg$ CSS $0.1493$ $0.1709$ $0.2039$ $0.2595$ $0.1293$ $0.1635$ $0.0214$ $0.0205$ Label Shift $0.2425$ $0.2425$ $0.2982$ $0.2982$ $0.1377$ $0.1377$ $0.0341$ $0.0348$
@@
1950-2007 1950-2009 1950-2011 DBLP/ACM Domains $2014-2016$ $2016-2018$ $2014-2016$ $2016-2018$ $2014-2016$ $2016-2018$ $D\rightarrow A$ $A\rightarrow D$ CSS $0.3005$ $0.3450$ $0.2583$ $0.3120$ $0.1833$ $0.2567$ $0.1573$ $0.2227$ Label Shift $0.2938$ $0.4396$ $0.2990$ $0.4552$ $0.2853$ $0.4438$ $0.3434$ $0.3435$
Appendix F Detailed Experimental Setup
F.1 Datasets Setup
For Arxiv, MAG, and DBLP/ACM we follow the setting of [59]. Hence we discuss the setting of synthetic dataset and pileup in the following: